# ASiM: Improving Transparency of SRAM-based Analog Compute-in-Memory Research with an Open-Source Simulation Framework

### v2

@@ Abstract
SRAM-based Analog Compute-in-Memory (ACiM) demonstrates promising energy efficiency for deep neural network (DNN) processing. Although recent aggressive design strategies have led to successive improvements on efficiency, there is limited discussion regarding the accompanying inference accuracy challenges. Given the growing difficulty in validating ACiM circuits with full-scale DNNs, standardized modeling methodology and open-source inference simulator are urgently needed. This paper presents SRAM ACiM Simulator, dubbed ASiM, a simulation framework specifically designed to assess inference quality, enabling comparisons of ACiM prototype chips and guiding design decisions. ASiM works as a plug-and-play tool that integrates seamlessly with the PyTorch ecosystem, offering speed and ease of use. Using ASiM, we conducted a comprehensive analysis of how various design factors impact DNN inference. We observed that activation encoding can tolerate certain levels of quantization noise, indicating a substantial potential for bit-parallel scheme to enhance energy efficiency. However, inference accuracy is susceptible to noise, as ACiM circuits typically use limited ADC dynamic range, making even small errors down to 1 LSB significantly deteriorates accuracy. This underscores the need for high design standards, especially for complex DNN models and challenging tasks. In response to these findings, we propose two solutions: Hybrid Compute-in-Memory architecture and majority voting to secure accurate computation of MSB cycles. These approaches improve inference quality while maintaining ACiM’s energy efficiency benefits, offering promising pathways toward reliable ACiM deployment in real-world applications. Code is available at https://github.com/Keio-CSG/ASiM
@@ I Introduction
Deep learning has brought revolutionary changes to a wide range of fields, not only in computer vision and natural language processing but also complex tasks spanning different modalities, delivering state-of-the-art performance. Recent advances have shown a scaling law in deep learning models, where performance consistently improves as the size of the model increases [1]. However, executing these large-scale models in traditional von Neumann architectures leads to substantial energy consumption and latency [2], which pose critical challenges, particularly for edge devices with limited resources. As a solution, Compute-in-Memory (CiM) has emerged as a novel paradigm aimed at overcoming these limitations by performing computations directly within the memory where data are stored [3]. This approach contrasts with conventional architectures that separate computation and memory units, CiM integrates computational capabilities into the memory itself, reducing data movement and thus substantially lowering energy consumption and latency.
Among the various CiM architectures, Analog CiM (ACiM) emerges as a promising approach to accelerate deep neural network (DNN) computations. Unlike Digital CiM (DCiM), which integrates digital logic within memory arrays, ACiM leverages the intrinsic analog characteristics of memory cells to perform multiply-and-accumulate (MAC) operations in the analog domain [4]. This approach inherently offers higher computational density and energy efficiency. Notably, SRAM-based ACiM has garnered significant attention due to its compatibility with existing CMOS processes. By incorporating capacitors into the SRAM array, the charge-domain SRAM-based CiM achieves high reliability while maintaining impressive energy efficiency [5], inspiring researchers to further push the limits of its performance. In recent years, state-of-the-art ACiM designs have tended to compress input activations and perform MAC operations in a bit-parallel fashion while lowering ADC bit precision to achieve improved energy efficiency [6, 7, 8, 9, 10].
@@
Despite of successive improvement on energy efficiency of ACiM designs, discussion regarding inference accuracy remains limited. This gap of understanding inference accuracy is particularly concerning by the challenges of validating ACiM circuits, because the massive scale of modern DNNs makes full hardware validation increasingly difficult. A more practical approach is partial validation, where researchers measure comprehensive characteristics of ACiM circuits and integrate them into software simulations to evaluate the end-to-end inference accuracy [11, 12, 6, 7]. Although this approach theoretically yields simulation results comparable to full hardware measurements, modeling ACiM circuits presents numerous latent challenges that can lead to misleading accuracy assessments. Moreover, the absence of standardized open-source ACiM simulators hampers reproducing circuit performance and making meaningful comparison across various research groups.
In this paper, we introduce SRAM ACiM Simulator, dubbed ASiM, an open-source simulation framework for designers to evaluate and compare the inference performance of their ACiM circuits, serving as a tool to guide design decisions and advance ACiM development. ASiM provides a comprehensive simulation environment that strictly follows SRAM-based ACiM computing rules while incorporating a variety of design factors to assess their effects on DNN inference. Unlike other frameworks, ASiM is a plug-and-play tool that integrates seamlessly into the widely used PyTorch ecosystem, offering speed and ease of use. Moreover, ASiM supports recent design innovations, such as bit-parallel operation, and extends its functionality to attention-based networks like Transformers. Through ASiM, we performed an in-depth analysis and numerical characterization of how various design factors affect inference accuracy in ACiM circuits. The main findings are as follows:
@@
  • Noise effects of ACiM circuits can severely impact inference accuracy due to the limited ADC dynamic range utilization in practical applications. Our analysis reveals that even minor ADC readout errors of 1 LSB can propagate through the network and cause significant accuracy degradation.
  • Inference on more complex tasks, such as ImageNet, is inherently more challenging than on CIFAR-10, requiring higher design standards or adoption of more reliable Hybrid CiM (HCiM) approach.
  • While bit-parallel ACiM can tolerate some quantization noise and thus significantly boost ACiM energy efficiency, it is more sensitive to analog noise, necessitating stricter design standards for ACiM circuits.
@@
  • We developed ASiM, an open-source framework for systematic ACiM accuracy evaluation that integrates with PyTorch and supports modern DNN architectures including Transformers.
  • With ASiM, we investigated the impact of key design factors on inference accuracy, and analyze the mechanism of how these factors degrade inference, revealing noise sensitivity issues that are usually overlooked before.
  • Building on our findings, we propose practical design guidelines and solutions to improve ACiM inference accuracy, enhancing its practicality for real-world applications.
The remainder of this paper is organized as follows. Section II reviews related works and provides preliminary concepts on ACiM and inference evaluation metrics. Section III describes the detailed implementation and modeling methodology of the ASiM framework. Section IV presents a comprehensive analysis of how each ACiM design factor impacts DNN inference accuracy and highlights fundamental limitations. Section V proposes simple yet effective solutions to improve ACiM inference accuracy. Finally, Section VII concludes the paper.
@@ II-A Charge Domain ACiM using SRAM
$O=\mathbf{x}\cdot\mathbf{w}=\sum_{p=0}^{P-1}\sum_{q=0}^{Q-1}(-1)^{k}{2}^{p+q}\sum_{n=0}^{N-1}x_{n}[p]w_{n}[q]$ (1)
@@
However, compressing more activation bits into a single ADC readout can potentially degrade inference accuracy, as it introduces higher quantization noise than the traditional bit-serial scheme. Additionally, recent designs often reduce ADC precision to achieve maximum energy efficiency, further amplifying quantization noise, which may compromise DNN inference. Under these conditions, the intrinsic non-idealities of analog computations in ACiM, such as noise and signal distortion, become increasingly problematic. This creates additional complexities in determining optimal design policies. Despite successive advancements in ACiM architectures, there remains a lack of comprehensive analysis on how these factors affect inference accuracy. To fill this critical gap, we have developed ASiM, a comprehensive simulation framework, to analyze these effects and highlight the bottlenecks in charge domain ACiM, providing insights for future ACiM design improvements.
@@ II-B Simulation Frameworks for ACiM
Validating a full DNN inference on a prototype ACiM chip is a challenging task that demands substantial effort. As a result, numerous simulation frameworks have been developed to model ACiM performance across various memory types. These frameworks not only evaluate the performance of fabricated chips but also assist in guiding chip specification decisions before ACiM circuit design begins [14, 15, 16, 17, 18]. However, these frameworks are primarily focused on circuit-level metrics, such as area and power consumption, and offer limited capabilities for evaluating inference accuracy on complete DNNs. Although there are simulation frameworks designed to assess DNN accuracy under non-ideal conditions [19, 20, 21], they primarily target non-volatile memory technologies, such as RRAM, which have different operational mechanisms and noise sources compared to SRAM. In RRAM-based ACiM, each memory cell is programmed to store multi-bit model weights, and multi-bit MAC operations are typically summed via current and measured by a single ADC readout. Consequently, noise in RRAM-based ACiM primarily arises from programming disturbances, resulting in weight variability. In contrast, SRAM-based ACiM leverages robust capacitors for near-deterministic accumulation, placing emphasis on the impact of ADC readout behaviors, such as thermal noise and random comparator decisions. As illustrated in Fig. 1, SRAM cells store single bits, requiring multi-bit weights to be decomposed into binary cycles, with results integrated through shift-and-add operations, adding complexity to the modeling process. In this paper, we demonstrate that although each ACiM cycle exhibits similar noise characteristics, their cumulative impact on final inference accuracy varies significantly. These factors necessitate careful consideration in constructing a reliable simulation framework for SRAM-based ACiM, particularly when evaluating inference accuracy on DNN models.
Saikia [22] introduces a modeling approach for inference accuracy in charge-domain SRAM ACiM. However, this model is restricted to a fixed configuration and supports only the ResNet model, limiting its applicability to modern ACiM designs and diverse DNN architectures. Furthermore, the approach uses a specialized setting that clips the ADC quantization range to a narrow range, which is challenging to implement in actual ADCs. As a result, the simulation outputs tend to overestimate inference accuracy, offering limited insights for ACiM design improvements.
In this work, we present a user-friendly simulation framework based on PyTorch, designed to be both general and practical for ACiM designers. By replacing PyTorch modules with ASiM modules, users can rapidly simulate predefined circuit metrics across various DNN models. Unlike most simulators created for RRAM, where each crossbar weight is modeled as a multi-bit value [21, 23], our ASiM framework aligns with SRAM CiM principles by decomposing computations into binary cycles and analyzing the influence of various design parameters on these MAC values, ensuring compatibility with fabricated circuits. By closely aligning with the computing operations of SRAM-based ACiM, this approach numerically exposes the varying impact of noise across different cycle indices. Additionally, ASiM integrates recent advancements in ACiM architecture and machine learning, including bit-parallel operations and support for Transformer models. This enhances its adaptability to a wide range of ACiM hardware configurations and enables detailed performance analysis across different models and tasks.
@@ II-C Evaluation Metrics for Inference Accuracy
To quantitatively assess the accuracy loss when using an ACiM accelerator for DNN inference, rigorous theoretical analysis and evaluation metrics have been established to guide the circuit design [24]. Among these metrics, the signal-to-quantization noise ratio (SQNR) and compute-signal-to-noise ratio (CSNR) are commonly used to characterize the inference quality of an ACiM chip. SQNR measures the signal-to-noise ratio (SNR) by considering only quantization effects during ACiM processing, while CSNR extends this by factoring in all non-idealities introduced by analog computations. The original definitions of SQNR and CSNR are given in [25] as:
$\text{SQNR}=\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{i}}}+\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{o}}}$ (2)
$\text{CSNR}=\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{i}}}+\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{o}}}+\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{\eta}}$ (3)
where ${\sigma}^{2}_{y}$ is the variance of the ideal floating-point output, and ${\sigma}^{2}_{y_{i}}$ and ${\sigma}^{2}_{y_{o}}$ represent the variance of the output considering input and output quantization, respectively. For CSNR, the additional term ${\sigma}^{2}_{\eta}$ accounts for variations caused by analog non-idealities. Jia [26] used randomly generated weight and activation vectors, comparing the post-quantization output to the ideal full-precision output to measure the SQNR of a real device, as described by:
@@
Although these results suggest a strong CSNR despite the presence of analog noise, our comprehensive investigation indicates that these evaluations may be overly optimistic. In particular, when CSNR is directly used to evaluate inference quality in software simulations, the resulting accuracy often appears unrealistically high. This issue stems from assumption that the vectors in the CSNR calculation are uniformly distributed, which does not align with the actual distributions found in DNN applications. While the noise terms in the denominator of Eqs. 2 and 3 can be minimized with careful circuit design, the signal term in the numerator is highly dependent on the specific DNN model and input data. Assuming a uniform distribution across the ADC’s full dynamic range can result in amplified signal variance, leading to an overestimation of CSNR.
In this paper, we demonstrate that the output distribution in ACiM depends on factors such as the DNN model, input activations, and MAC cycle index, with the real output signal typically confined to a smaller range. As a result, the CSNR metric may not effectively capture inference accuracy in real DNN applications. Our ASiM framework, therefore, serves as a more accurate tool for evaluating inference performance in practical DNN tasks and supports design space exploration for ACiM chip design.
@@ III-A Overview
[…]
@@ III-B1 Algorithm
2:  Quantize input activations and weights to $X$ and $W$ bits. Convert them into binary bit tensors and split them into $N$ tensors to match the dimensions of the macros
@@ III-B5 Analog Noise Modeling
As detailed in Algorithm 1, analog noise can be added to the CBL voltage of each column to emulate real ACiM devices, enabling evaluation of its effect on inference accuracy. ASiM includes two noise models: random noise and non-linearity, both of which can be configured with user-defined noise intensities. The combined analog noise model, incorporating these two noise types, reflects the majority of conditions found in fabricated circuits. Static offset noise, while potentially present in real circuits, can be mitigated through ADC code calibration and is therefore omitted from ASiM for simplicity.
Random noise covers sources such as $kT/C$ noise and random comparator decision errors, which can be simply represented with a Gaussian distribution modeling an ADC input-referred noise [28, 29]. The distribution has a mean of zero and a standard deviation corresponding to the noise intensity, with ASiM taking this intensity as a percentage relative to the entire operational voltage range. Non-linearity arises from device mismatches and layout-specific issues; ASiM incorporates capacitance mismatch in local memory cells and DACs, scaled by $\sqrt{N}$, where $N$ is the number of capacitors in the CBL and DAC [30]. Users can also incorporate custom noise models to account for non-linearities specific to their designs, enabling tailored inference performance analysis with ASiM.

### v3

# ASiM: Modeling and Analyzing Inference Accuracy of SRAM-Based Analog CiM Circuits

@@ Abstract
SRAM-based Analog Compute-in-Memory (ACiM) demonstrates promising energy efficiency for deep neural network (DNN) processing. Nevertheless, efforts to optimize efficiency frequently compromise accuracy, and this trade-off remains insufficiently studied due to the difficulty of performing full-system validation. Specifically, existing simulation tools rarely target SRAM-based ACiM and exhibit inconsistent accuracy predictions, highlighting the need for a standardized, SRAM CiM circuit-aware evaluation methodology. This paper presents ASiM, a simulation framework for evaluating inference accuracy in SRAM-based ACiM systems. ASiM captures critical effects in SRAM based analog compute in memory systems, such as ADC quantization, bit parallel encoding, and analog noise, which must be modeled with high fidelity due to their distinct behavior in charge domain architectures compared to other memory technologies. ASiM supports a wide range of modern DNN workloads, including CNN and Transformer-based models such as ViT, and scales to large-scale tasks like ImageNet classification. Our results indicate that bit-parallel encoding can improve energy efficiency with only modest accuracy degradation; however, even 1 LSB of analog noise can significantly impair inference performance, particularly in complex tasks such as ImageNet. To address this, we explore hybrid analog–digital execution and majority voting schemes, both of which enhance robustness without negating energy savings. ASiM bridges the gap between hardware design and inference performance, offering actionable insights for energy-efficient, high-accuracy ACiM deployment. Code is available at https://github.com/Keio-CSG/ASiM
@@ I Introduction
Deep learning has brought revolutionary changes to a wide range of fields, not only in computer vision and natural language processing but also complex tasks spanning different modalities, delivering state-of-the-art performance. Recent advances have shown a scaling law in deep learning models, where performance consistently improves as the size of the model increases [1]. However, executing these large-scale models in traditional von Neumann architectures leads to substantial energy consumption and latency [2], which pose critical challenges, particularly for edge devices with limited resources. Compute-in-Memory (CiM) has emerged as a novel paradigm to overcome these limitations by performing computations directly within the memory arrays  [3]. By reducing data movement between logic and memory, CiM architectures can achieve significant improvements in both energy efficiency and throughput.
Among CiM architectures, analog compute-in-memory (ACiM) offers high computational density and energy efficiency by leveraging the analog properties of memory cells to perform multiply-and-accumulate (MAC) operations directly in the analog domain [4]. SRAM-based ACiM, in particular, has attracted attention due to its compatibility with standard CMOS processes and reliable charge-domain operation [5]. In recent years, state-of-the-art ACiM designs have tended to compress input activations and perform MAC operations in a bit-parallel fashion while lowering ADC bit precision to achieve improved energy efficiency [6, 7, 8, 9, 10]. These techniques, while beneficial in some respects, exacerbate quantization errors and make it difficult to preserve inference accuracy, highlighting the importance of accurate and thorough evaluation.
@@
Despite of successive improvement on energy efficiency of ACiM designs, discussion regarding inference accuracy remains limited. This gap of understanding inference accuracy is particularly concerning by the challenges of validating ACiM circuits, because the massive scale of modern DNNs makes full hardware validation increasingly difficult. A more practical approach is partial validation, where researchers measure comprehensive characteristics of ACiM circuits and integrate them into software simulations to evaluate the end-to-end inference accuracy [11, 12, 6, 7]. Although this approach theoretically yields simulation results comparable to full hardware measurements, modeling ACiM circuits presents numerous latent challenges that can lead to misleading accuracy assessments.
We examined several existing simulation frameworks aimed at assessing inference accuracy and observed significant variation in their predictions. These discrepancies suggest that current tools may struggle to provide consistent guidance for design validation or silicon correlation. A key contributing factor is the lack of standardized, open-source simulators tailored specifically to SRAM-based ACiM. The inconsistent predictions produced by different tools highlight a key research question: What modeling components are essential for accurate inference evaluation in SRAM-based ACiM systems, and how can we ensure reliable, consistent predictions across diverse design scenarios? Addressing this gap would support the development of standardized evaluation practices and facilitate the broader adoption of ACiM architectures.
In this paper, we introduce ASiM, an accurate simulation framework that bridges low-level circuit behavior and high-level inference performance in SRAM-based ACiM systems. ASiM models essential circuit effects such as ADC quantization, bit-parallel encoding, and analog noise, enabling accurate evaluation of how hardware design choices impact DNN inference accuracy. Designed for practical use, ASiM integrates seamlessly with PyTorch as a plug-and-play tool, providing both modeling fidelity and ease of deployment. It supports SRAM ACiM specific architectural innovations, including bit-parallel execution, and is compatible with modern DNNs such as Transformers. Using ASiM, we conduct a detailed analysis of how various circuit-level factors influence inference accuracy. The main findings are as follows:
@@
  • Analog noise within ACiM circuits severely impact inference accuracy: Our analysis reveals that even minor ADC readout errors of 1 LSB can propagate through the network and cause significant accuracy degradation.
  • Complex tasks demand stricter hardware design constraints: Inference on tasks like ImageNet is substantially more sensitive to circuit-level nonidealities than simpler benchmarks such as CIFAR-10. Maintaining accuracy on such tasks requires tighter design margins or the adoption of more robust approaches like Hybrid CiM (HCiM).
  • Bit-parallel ACiM improves energy efficiency but increases sensitivity to noise: While this architecture tolerates moderate quantization noise and achieves substantial energy savings, it is more susceptible to analog noise, necessitating stricter design standards to ensure inference reliability.
@@
  • We develop ASiM, an open-source simulation framework for systematic inference accuracy evaluation in SRAM-based ACiM systems. ASiM integrates with PyTorch and supports modern DNN architectures, including Transformers.
  • With ASiM, we analyze the impact of key circuit-level design factors on inference accuracy and uncover how specific noise sources degrade model performance, revealing sensitivity issues often overlooked in prior work.
  • Building on our findings, we provide practical design guidelines and solutions to improve ACiM inference accuracy, enhancing its practicality for real-world applications.
The remainder of this paper is organized as follows. Section II reviews related works and provides preliminary concepts on ACiM and inference evaluation metrics. Section III describes the detailed implementation and modeling methodology of the ASiM framework. Section IV presents a comprehensive analysis of how each ACiM design factor impacts DNN inference accuracy and highlights fundamental limitations. Section V provides simple yet effective solutions to improve ACiM inference accuracy. Finally, Section VII concludes the paper.
@@ II-A Charge Domain ACiM using SRAM
$O=\mathbf{x}\cdot\mathbf{w}=\sum_{p=0}^{P-1}\sum_{q=0}^{Q-1}(-1)^{k}{2}^{p+q}\sum_{n=0}^{N-1}x_{n}[p]w_{n}[q].$ (1)
@@
However, compressing more activation bits into a single ADC readout can potentially degrade inference accuracy, as it introduces higher quantization noise than the traditional bit-serial scheme. Additionally, recent designs often reduce ADC precision to achieve maximum energy efficiency, further amplifying quantization noise, which may compromise DNN inference. Under these conditions, the intrinsic nonidealities of analog computations in ACiM, such as noise and signal distortion, become increasingly problematic. This creates additional complexities in determining optimal design policies. Despite successive advancements in ACiM architectures, there remains a lack of comprehensive analysis on how these factors affect inference accuracy. To fill this critical gap, we have developed ASiM, a comprehensive simulation framework, to analyze these effects and highlight the bottlenecks in charge domain ACiM, providing insights for future ACiM design improvements.
@@ II-B Simulation Frameworks for ACiM
Full DNN inference validation on ACiM prototypes requires significant resources. As an alternative, various simulation frameworks have been developed to model ACiM performance across memory types. These tools aid both pre-silicon design exploration and post-silicon performance evaluation. Some emphasize circuit-level performance modeling [14, 15, 16], while others incorporate analog nonidealities to assess inference accuracy [17, 18, 19, 20]. Frameworks such as X-PIM [21] aim to address both.
However, most existing frameworks are optimized for non-volatile memories like RRAM and Flash, whose noise characteristics differ from SRAM. In contrast, ASiM adheres to SRAM-specific modeling principles, using binary cycle decomposition and fine-grained modeling of MAC outputs to reflect real fabricated circuit behavior. Implemented within PyTorch, ASiM supports recent ACiM features including bit-parallel operations and Transformer-based models, enabling flexible, accurate performance evaluation across diverse workloads.
@@ II-C Evaluation Metrics for Inference Accuracy
To quantify inference degradation from ACiM execution, metrics such as signal-to-quantization noise ratio (SQNR) and compute-signal-to-noise ratio (CSNR) are commonly used [22]. SQNR captures noise from quantization alone, while CSNR accounts for additional analog effects. The original definitions of SQNR and CSNR are given in [23] as:
$\text{SQNR}=\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{i}}}+\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{o}}},$ (2)
$\text{CSNR}=\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{i}}}+\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{y_{o}}}+\frac{{\sigma}^{2}_{y}}{{\sigma}^{2}_{\eta}},$ (3)
where ${\sigma}^{2}_{y}$ is the variance of the ideal floating-point output, and ${\sigma}^{2}_{y_{i}}$ and ${\sigma}^{2}_{y_{o}}$ represent the variance of the output considering input and output quantization, respectively. For CSNR, the additional term ${\sigma}^{2}_{\eta}$ accounts for variations caused by analog nonidealities. Jia [24] used randomly generated weight and activation vectors, comparing the post-quantization output to the ideal full-precision output to measure the SQNR of a real device, as described by:
@@
While prior studies report high CSNR despite analog noise, our findings indicate these values may not generalize well. Many such estimates assume uniformly distributed activation and weight vectors, leading to inflated signal variance in the numerator of Eq. 3. In actual DNN applications, outputs tend to occupy narrower ranges, reducing true signal power and distorting CSNR interpretation.
In this paper, we demonstrate that the output distribution in ACiM depends on factors such as the DNN model, input activations, and MAC cycle index. Consequently, CSNR may not reliably reflect inference quality in practice. Our ASiM framework offers a more representative evaluation for realistic DNN tasks and supports design space exploration grounded in accurate modeling.
@@ III-A Overview
ASiM is implemented within PyTorch to ensure seamless integration with modern DNN workflows. It provides a plug-and-play simulation framework tailored for SRAM-based charge-domain ACiM. For MAC-heavy operations such as CONV, Linear, and Attention, ASiM modules can replace standard PyTorch layers, enabling bit-wise simulations with user-defined circuit parameters.
During inference, quantized weights and activations are decomposed into binary bit tensors, split according to the predefined macro dimensions, and mapped onto the ACiM macros. The bit-wise simulation transforms the original floating-point operations into $\text{wbit}\times\text{xbit}\times\text{mapping}$ operation is simulated using native PyTorch functions, producing ideal analog voltages on CBLs. ASiM then applies analog noise and ADC effects to these voltages. By leveraging PyTorch’s CUDA acceleration, ASiM achieves both simulation fidelity and high performance. Additionally, ASiM supports bit-parallel operation, where input activations are encoded based on encoding bit width, and the framework calculates bit-parallel voltage values using pre-configured algorithms, ensuring compatibility with most latest ACiM designs.
@@ III-B1 Algorithm
2: Quantize input activations and weights to $X$ and $W$ bits. Convert them into binary bit tensors and split them into NN tensors to match the dimensions of the macros
@@ III-B5 Analog Noise Modeling
As detailed in Algorithm 1, analog noise can be added to the CBL voltage of each column to emulate real ACiM devices, enabling evaluation of its effect on inference accuracy. As ASiM is designed as a system-level simulator, it does not model or differentiate between specific low-level analog noise sources. Instead, the cumulative effect of these noise sources is abstracted as equivalent fluctuations on CBL voltage for ADC readout. Therefore, ASiM includes two simple noise models: random noise and nonlinearity, both of which can be configured with user-defined noise intensities. The combined analog noise model, incorporating these two noise types, reflects the majority of conditions found in fabricated circuits. Static offset noise, while potentially present in real circuits, can be mitigated through ADC code calibration and is therefore omitted from ASiM for simplicity.
Random noise covers sources such as $kT/C$ noise and random comparator decision errors, which can be simply represented with a Gaussian distribution modeling an ADC input-referred noise [26, 27]. The distribution has a mean of zero and a standard deviation corresponding to the noise intensity, with ASiM taking this intensity as a percentage relative to the entire operational voltage range. Nonlinearity arises from device mismatches and layout-specific issues; ASiM incorporates capacitance mismatch in local memory cells and DACs, scaled by $\sqrt{N}$, where $N$ is the number of capacitors in the CBL and DAC [28]. Due to the stochastic nature of mapping DNNs to ACiM hardware, spatial noise such as mismatch and PVT variation, is expected to have a similar impact on accuracy and is thus integrated into the overall random noise and nonlinearity model in this paper. Users can also incorporate custom noise models to account for nonlinearities specific to their designs, enabling tailored inference performance analysis with ASiM.